OpenStudy (anonymous):

-3/4n=-36 what does n equal pleas help if you know.

OpenStudy (wolf1728):

When we have an equation sch as a/b = c/d we can cross-multiply to get a*d = b*c Do you think you could do that with the -3/4n = -36 equation?

jhonyy9 (jhonyy9):

so first of all this -3/4n wan being -3/(4n) or (-3n)/4 ? - i think the second so than - because on both sides there are minus ,first you divide both sides by -1 and will get 3 ---n = 36 4 3n=4*36 - divide both sides by 3 n= 4*12 n= 48